How can I make it easier for myself to understand steps to solving rational equations like this one... x-5/x-8 = x-1/x-5
Answer by checkley77(12844)

(x-5)/(x-8)=(x-1)/(x-5) if this is what you mean THEN:
Cross multiply.
(x-8)(x-1)=(x-5)(x-5)
x^2-9x+8=x^2-10x+25 cancel out thw x^2 terms & combine like terms.
-9x+10x=25-8
x=17 ans.
Proof:
(17-5)/(17-8)=(17-1)/(17-5)
12/9=16/12
9*16=12*12
144=144